OpenStudy (anonymous):

x/5+6=9 help on solving equation?

OpenStudy (lgbasallote):

don't you mean 6?

OpenStudy (anonymous):

Im pretty sur but how do i solve for the fraction?

OpenStudy (anonymous):

Again typo!

OpenStudy (lgbasallote):

lol. darn i was about to point that out

OpenStudy (anonymous):

Go step by step, will you just show us how you move 6 to the right?

OpenStudy (anonymous):

When i move 6 to the right its x/5=15 ...

OpenStudy (anonymous):

I'm pretty freak out today that too many askers told me that they understand, but it turns out ... ( sighs :'( )

OpenStudy (anonymous):

Pls people just make life easier by admitting where you don't understand!

OpenStudy (anonymous):

I dont understand how to make it in x=# form. what is the next step after x/5=15? any better?

OpenStudy (anonymous):

@gabbbdee so you don't know how to move 6 By switching the side, you need to switch the SIGN: 6 -> -6, Pls learn the rule!

OpenStudy (anonymous):

no i know that rule, but after you move the number what is the next step?

OpenStudy (anonymous):

3**

OpenStudy (anonymous):

Then why you make it basic mistake ( sighs ) x/5 = 3 ->x = 3 * 5 = ...